[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

80.0. "ASHLAND CHEMICAL - SLURRY,AMMONIUM HYDROXIDE" by SUBPAC::BJUBINVILLE (CMP-6 Equipment Engineering) Wed Nov 20 1996 21:02

    
        This note is dedicated to ASHLAND CHEMICAL replies
      for passdowns, readings, filter changes ect...
T.RTitleUserPersonal
Name
DateLines
80.1Ashland Chemical contact numbersSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Nov 20 1996 21:066
    
       Ashland Chemical 24-hotline number   1-800-489-4393
    
       Dave Agresti - site manager  x6112
    
       Ashland office - 6122
80.210" filter installedSUBPAC::LANDRYMon Nov 25 1996 21:506
    Today, at approximately 12:00 PM (noon), the slurry filter was changed
    over to a 10" filter.  This is the first attempt with this filter. 
    Ashland will check on it more frequently than normal.  
    
                                            A shift
    
80.310" removedSTRATA::TDYERMon Dec 02 1996 07:377
    
    Ashland removed the 10" filter-rite 10" filter on Friday at 4:00pm,
    Osmonics put back on line. The filter was running fine, they replaced
    it because of fear it may clogg over the weekend. Testing will continue
    this week.
    
    Tim D.
80.4Title : Removed 10"HOTLNE::INGLESMon Dec 02 1996 07:494
The 10" filter in slurry ran monday - friday with no
problems. We switched back to the osmonics to prevent
from occuring over the weekend. The filter will be 
changed saturday and sunday.  
80.5Removed 10" / Changed Osomics FilterHOTLNE::OUELLETTEMon Dec 02 1996 08:417
The Osmonics filter in slurry was changed out Saturday November 30 1996.

Also Sunday November 31 1996 it was changed again.

Both filters were changed between the hours of three & four pm each day.

MAO
80.6millipore filterHOTLNE::SPANGTue Dec 03 1996 18:317
Today Tuesday 12/3/96 We installed the millipore filter housing and millipore
filter # cr2501006 a 25 um. It was turned on at 17:45 and the presure readings 
where 40 psi in and 38 psi out. The Osmonics filter is in place if need be due
to a failurem of the millipore.


Dave Spang 674 # 1 
80.7Filterite 20" carteridgeHOTLNE::OUELLETTETue Dec 17 1996 15:186
    12/16/96 at 4pm the Filterite 20" 30um filter was turned on-line.
    
    The dispense out line pressure was at 35psi directly after the filter
    was turned on-line.
    
    MAO/RM
80.8filter changeHOTLNE::SPANGFri Dec 27 1996 11:229
Today Friday Dec 27 We changed out the  filter-rite filter in the
 bulk slurry.system. We have switched back to the Osmonics filter
 at this point and will run on the Osmonics until Monday. We will 
change the Osmonics filter out on Saturday afternoon at 4:00 and put 
another Osmonics filter on line. On Monday we will switch to the 20"
Pall filter. 


				Dave Spang 
80.9nHOTLNE::INGLESMon Jan 06 1997 11:192
    On 1/3/97 at 0830 we switched from the Osmonics filters to the 20" Pall
    filter. We will run on this filter until it fails.
80.10Pall Filter DiedHOTLNE::OUELLETTEThu Jan 09 1997 10:062
    On 1/9/97 at 0830 we switched from the 20"pall filter to a 30um
    Osmonics. 
80.11pal failedSTRATA::TDYERThu Jan 09 1997 10:247
    
    The Pal ultra-pleat filter failed  for particles. B2 failed on 1/7
    and B1 on 1/9. This was the 2nd test for this filter. We are now
    confident that this filter is not acceptable. Evaluation will continue
    with the Filter-Rite and Milli-pore over the next few weeks.
    
    Tim 
80.12high particlesHOTLNE::INGLESThu Jan 09 1997 16:156
    At 1330 we drained the slurry system and flushed it with DI water for 
    30 minutes. The Osmonic filters were removed and new Osmonics filters were 
    installed. The system was filled with slurry and was turned over to CMP
    at 1500.
    
    		George of Ashland
80.13testing 20" Filterite CMP 630-20UV300.HOTLNE::INGLESFri Jan 10 1997 15:384
    At 1530 we switched to the 20" Filterite using filter# CMP 630-20UV300.
    We will run with this until failure. 
    
    		George Of Ashland
80.15filter changeHOTLNE::MULLINFri Jan 17 1997 10:578
On 1-16-97 we changed back to osmonic filter at 6:30am
due to system clog, after running filterite cartridge(20")
Found filter and housing to be caked with slurry sludge.

On 1-17-97 we changed over to the opposite osmonic filter due to
high pressure readings. This is the left side filter.

				RPM
80.16HOTLNE::OUELLETTEMon Jan 20 1997 11:285
    01/20/97 The left side Osmonics filter had extremely high pressure
    readings so the switch over to the right side osmonics was completed
    11:30am.
    
    MAO
80.17changed osmonics filter 30 micronHOTLNE::INGLESWed Jan 22 1997 00:263
    At 11:40 pm the slurry filter climbed to a 30 lb differential. We
    switched to the left osmonics filter and changed out the right. We are
    still using the 30 micron osmonics.
80.18changed osmonics filter 30 micronHOTLNE::MEHLERThu Jan 23 1997 21:282
    At 8:30 pm the slurry filter was switched to right osmonics and the
    left was changed out.  We are still using 30 micron osmonics.
80.19Megasys Coming Soon?HOTLNE::OUELLETTEFri Jan 24 1997 08:335
This morning the CMP 630-10UV-M3 300 Filterite filter was installed.

Tonight this filter will be turn on-line @9pm, unless the osmonics fails.

80.2010" filterite now operatingHOTLNE::MEHLERFri Jan 24 1997 22:522
    The 10" 30 micron filterite slurry filter began running at 10:30 pm
    
80.21osmonics filter back on lineHOTLNE::SPANGMon Jan 27 1997 14:005
    On Sunday jan 26 at 11:50 pm the 10" memtec slurry filter clogged
    We went back to the osmonics filter and will changes these out every 48
    hours untill further notice.
    
    				DS/GI
80.2280.22HOTLNE::MULLINMon Feb 03 1997 08:205
    Saturday, Feb 2, 4:30p, changed out left osmonic filter, also changed 
    out the right as well. Unit is now running on right, with left ready.
    
    Filter differential was 8 lbs.
    
80.23Updating Filter ChangesHOTLNE::OUELLETTEMon Feb 24 1997 17:386
    2/23/97 
    
    	Changed from the left side osmonics to the right side osmonics
    filter. The time was 6:30pm Sunday. 
    
    MAO
80.24HOTLNE::OUELLETTETue Feb 25 1997 20:195
2/25/97 7pm The right side osmonics filter was changed out to the left side.

A new outlet pressure gauge was installed to the right side filter housing.

mao
80.25HOTLNE::OUELLETTESun Mar 02 1997 19:457
    2/28/97 The right side osmonics was changed over to the left side
    filter housing. Roughly around 7pm.
    
    3/2/97 The left side osmonics was changed over to the right side
    filter housing @ 7pm.
    
    MAO
80.26Slurry filter changeHOTLNE::MEHLERWed Mar 12 1997 16:003
System chemistry osmonic slurry filter was changed last night
at 9:30 pm.  It was changed over to the right side.
RCM
80.27Osmonic filter changeHOTLNE::MEHLERThu Mar 13 1997 21:033
The Osmonic Slurry filter was changed at 8:30pm.  It was 
changed over to the left side.
RCM
80.28Slurry Osmonic filter changedHOTLNE::MEHLERMon Mar 17 1997 18:404
The Osmonic filter on the systems chemistry Slurry system
was changed at 5:30pm on 3-17-97.  It was switched over to
the left side.
-rcm
80.29slurry osmonic filter changeHOTLNE::INGLESSun Mar 23 1997 20:232
    The Osmonic filter was changed on 3/21/97 at 4pm. The system is running
    on the right osmonic filter.
80.30slurry osmonic filter changeHOTLNE::INGLESSun Mar 23 1997 20:252
    The Osmonic filter was changed on 3/23/97 at 4:30 pm. The system is
    running on the left osmonic filter.
80.31Osmonic filter changeHOTLNE::MEHLERTue Mar 25 1997 23:262
After notifying CMP, the Osmonic slurry filter was changed at approx. 4:30.  It 
was switched to the right side.
80.32Osmonic filter changeHOTLNE::MEHLERThu Mar 27 1997 19:373
After coordinating a time with CMP, the Slurry osmonic filter was changed tonight
at 6:00pm.  The filter was changed to the right side.  
-RCM and GI
80.33Slurry osmonic filter changeHOTLNE::MEHLERMon Mar 31 1997 20:432
The Slurry osmonic filter was changed tonight at 5:30pm.  It was changed to the
right side by RCM and GI.
80.34new slurry filter CMP630-20UV-M3 300HOTLNE::INGLESThu Apr 03 1997 00:165
    At 6:00 PM we changed the slurry system filter. We are now using the 20
    inch Ametec filter housing and filter # CMP630-20UV-M3  300. We have
    both Osmonic filter housings with new filters installed for backup.
    
    			GI and RCM
80.35HOTLNE::OUELLETTEFri Apr 04 1997 23:289

	Today @ 1120 am the slurry filter was changed with a new 
	cmp630-20uv-m3 300.

	MAO


	Today @ 415pm the Mega slurry system filter was changed for CMP.B4.
80.36HOTLNE::OUELLETTETue Apr 08 1997 17:139
    Today 4/8/97 the MegaflowII dispense filter was replaced with a new
    fitler cmp 630 20 m3 300
    
    Also the SC112 filter was replaced with a new filter
    cmp 630 20 m3 300.
    
    Both filters were changed at 11am.
    
    MAO
80.37Mega and Sys Chem slurry filter changesHOTLNE::INGLESFri Apr 11 1997 12:583
    Today we changed both the System Chemistry and the Mega CDU filters. We
    are using filter number CMP 630-20UV-M3  300. These filters were
    changed between 11:00am and 11:30am. 
80.38HOTLNE::OUELLETTETue Apr 15 1997 12:033
	Today we changed both the system chemistry and mega cdu filters.
	We are using cmp-20uv-m3 300. These filters were changed between 
	10:30 and 11am.
80.39HOTLNE::OUELLETTEFri Apr 25 1997 12:0110
	APRIL 18TH 1997

	The filter to system chemistry MDU was changed ~10:30am.

	The filters to Mega MDU was changed ~10:30am.

	All filters replaced with cmp-20uv-m3 300 type filters.

	
80.40HOTLNE::OUELLETTEFri Apr 25 1997 12:033
	Today at 1015am the slurry filter was changed using a cmp-20uv-m3 300
	type filter.
80.41Mega Distribution 2 slurry filter change.HOTLNE::INGLESMon Apr 28 1997 12:355
    Today at 1000 we changed the Distribution 2 slurry filter using
    cmp-20uv-m3  300. We did this in order to repair the flter housing 1/4
    turn valve.
    
    			GI
80.42System Chemistry and Mega Slurry filter ChangeHOTLNE::INGLESTue Apr 29 1997 11:212
    Today at 0945 we changed the system chemistry and Mega ditribution one 
    filters using CMP-20UV-M3  300. 
80.43System Chemistry and Mega System slurry filter Change.HOTLNE::INGLESFri May 02 1997 14:402
    Toaday we changed filters on the System Chemistry and Distro. 1 and 2 on
    the Mega System using the CMP-20UV-M3  300 filters.
80.44HOTLNE::OUELLETTETue May 06 1997 14:102
Today we changed filters on the system chemistry and distro 1 & 2 on 
the Mega System using CMP-20UV-m3 filters.
80.45HOTLNE::OUELLETTEFri May 09 1997 12:072
Today May 9, 1997 bulkchemical group changed filters to SCI & MEGAFLOWII
distrobution dispense systems using cmp-200uv-m3.
80.46HOTLNE::MULLINTue May 13 1997 11:412
    5/13/97 - Ashland changed out filters in both the SCI & MEGASYS. slurry 
    distrubution units 10am-1030
80.47ASHLAND CHEMICAL - SLURRY FILTER CHANGEHOTLNE::MEHLERTue May 13 1997 21:443
5/13/97- At the request of CMP the system's chemistry filter was changed a 
second time due to very high particle counts. 8:30-8:40pm
-rcm
80.48Systems Chemistry Slurry System filter changeHOTLNE::MEHLERFri May 16 1997 04:365
5/16/97
	On call responded to CMP's page that slurry from the systems chemistry
CDU was not reaching CMP.  Found that filter was 100% clogged and changed 
immediately.  Finished at 3:00am.
-rcm
80.49Slurry filter changeHOTLNE::MEHLERSat May 17 1997 16:423
    5/17/97
    At the request of CMP and due to high particles the Mega System Slurry
    filters were changed at 3:00pm by RCM.
80.50Slurry filter changesHOTLNE::INGLESMon May 19 1997 14:554
    Today at 0800 the Slurry Mega system filters Dist.1&2 were changed. At 1330
    the Slurry System Chemistry filter was changed.
    
    				GI
80.51HOTLNE::OUELLETTEWed May 21 1997 16:105
Today at 10:15am the System Chemistry distrobution filter was changed.

Also at 2pm the MegaFlowII distrobution filters were changed.

MAO 
80.52NOTE 80.52HOTLNE::MULLINSat May 24 1997 14:154
    This is a late entry to notify all the changeout of pump # 1 in the
    Sys. Chem. CDU on Thursday, 5/22/97. There was zero downtime to the
    FAB.
    
80.53HOTLNE::OUELLETTETue May 27 1997 08:211
Friday May 23, the System Chemistry CDU filter was changed.
80.54Slurry System Filter ChangesHOTLNE::INGLESWed May 28 1997 09:412
    At 8:00 am the System Chemistry and Mega System Slurry filters were
    changed. 
80.55Slurry System Filter ChangesHOTLNE::INGLESMon Jun 02 1997 16:152
    At 8:00 am the System Chemistry Slurry filter was changed. At 2:30 pm
    the Mega System Slurry filter was changed.
80.56Slurry System Filter ChangesHOTLNE::INGLESWed Jun 04 1997 11:342
    At 0900 we changed the System Chemistry Slurry filter and at 0945 we
    changed the Mega System Slurry filters.